OpenStudy (anonymous):

How do you evaluate the integral using the substitution method? x/(4x^2 - 3)^0.4

OpenStudy (anonymous):

u is the expression in the parenthesis of the denominator

OpenStudy (anonymous):

so du=8x dx

OpenStudy (anonymous):

then make an integral with u's in it

OpenStudy (anonymous):

\[\int\frac{x}{(4x^2-3)^{0.4}}dx\] say u=(4x^2-3) then du/dx=8x or xdx=du/8 the integral becomes: \[\int\frac{1}{8(u)^{0.4}}du\]

OpenStudy (anonymous):

So you just put 8x in for u?

OpenStudy (anonymous):

No, @imron07 set it up perfectly... solve the u integral... then put the x's back in after taking the integral by substituting u=4x^2-3 where ever u is.

OpenStudy (anonymous):

you could see it as this\[\frac{1}{8}\int\limits_{}^{}u ^{-0.4}du\]use the power rule for integrals after you are done, put the x's back in

OpenStudy (anonymous):

Oh ok thank you guys!

OpenStudy (anonymous):

No problem :)